What percent by mass of bromine does the insectici
Answer #1

24) C2H4Br2 has 2C + 4H + 2Br , so total molar mass = 2 x 12 + 4 x 1 + 2 ( 80) = 188

Br mass % = ( mass of Br / molar mass of compound) = 100 x ( 2 x 80 /188) = 85.1 %

25) Molecular formula = emperical formula x number of repeating units = ( CHCl) x n

emperical formula mass = ( 12 + 1 + 35.5) = 48.5 amu

molar mass = 48.5 x n = 290.8

n = 6

hence Molecular formula is ( CHCl)6 i.e C6H6Cl6